OpenStudy (anonymous):

7/1x > - 1/4 use multiplaction graph solution

OpenStudy (helder_edwin):

\[ \large \frac{7}{x}>-\frac{1}{4} \] ??

OpenStudy (anonymous):

it says use mulitplication to sove inequality and grpah the solution

OpenStudy (helder_edwin):

but. is this the inequation?

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

OK \[ \large \frac{7}{x}>-\frac{1}{4} \] \[ \large \frac{7}{x}+\frac{1}{4}>0 \] \[ \large \frac{28+x}{4x}>0 \] following??

OpenStudy (anonymous):

not at all :(

OpenStudy (helder_edwin):

which part?

OpenStudy (anonymous):

i got -1/14

OpenStudy (anonymous):

but having trouble how to grpah in a line grpah

OpenStudy (helder_edwin):

wait until i finish. OK?

OpenStudy (helder_edwin):

u solve \[ \large 28+x=0\Rightarrow x=-28 \] \[ \large 4x=0\Rightarrow x=0 \] this divides the (real) line into three parts

OpenStudy (helder_edwin):